OpenStudy (anonymous):

175 mL of methanol (CH4O) is mixed with 75 mL of water. what is the concentration in moles/liter of water in this solution? Density H2O = 1.00 g/mL

OpenStudy (aaronq):

in this case water is the solute. you need to know the mass of the water, typically you'd need to use the density but since it's 1 g/mL, it's the same thing. so 75 mL of water = 75 g of water. now find the moles: \(n=\dfrac{m}{M}\) then find the molarity \(M=\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (anonymous):

ok but how do i do that?

OpenStudy (aaronq):

use the formulas i posted

OpenStudy (anonymous):

hmm tbh i have no idea

OpenStudy (aaronq):

do you know what the symbols, "n, m, M" mean?

OpenStudy (anonymous):

no ;/ i just started the class so and have a hard time getting whats going on:(

OpenStudy (aaronq):

hm it seems like you need to read your book.. n=moles m=mass M=molar mass (in the first equation), and Molarity (in the second equation) Do you know how to find the molar mass of water?

OpenStudy (anonymous):

yeah H2O - made up of 2 Hydrogen atoms and 1 Oxygen atom. Mass of 1 Hydrogen = 1g Mass of 1 Oxygen = 16g 1x2 = 2+16 = 18g

OpenStudy (anonymous):

you have to find the molar mass of each in the H2O the add them to get the molar mass

OpenStudy (aaronq):

okay, so after you find the moles using the first equation: \(n_{H_2O}=\dfrac{75g}{18g/mol}\)

OpenStudy (anonymous):

so then you'd get 4.16 moles?

OpenStudy (aaronq):

okay, now you use the formula \(Molarity=\dfrac{n_{H_2O}}{L_{solution}}\) how many liters are there in solution?

OpenStudy (aaronq):

let me rephrase that..how many liters make up the solution?

OpenStudy (anonymous):

0.075 L ? bc you have to convert mL to L

OpenStudy (aaronq):

yeah, thats how much theres is of water, but you're forgetting about the methanol

OpenStudy (anonymous):

the methanol would be 0.175 L

OpenStudy (aaronq):

so the volume of the solution is?

OpenStudy (anonymous):

would i use the methanol or water to find the volume?

OpenStudy (aaronq):

you would use both, the volume (in L) of the solution is the space the solution occupies. \(L_{solution}=L_{methanol}+L_{water}\) btw, a solution is all of the components. the solute is what there is LESS of the solvent is what there is MORE of usually water is the solvent, but they reversed in this case.

OpenStudy (anonymous):

so it'd = 0.25

OpenStudy (anonymous):

0.175+0.075 = 0.25

OpenStudy (aaronq):

yeah thats right. now plugging into the formula: \(Molarity=\dfrac{4.16 \;moles}{0.25\; L}\)

OpenStudy (anonymous):

= 16.64 moles L

OpenStudy (aaronq):

it should be "moles/L" which is what molarity is. don't get confused about further problems where water is typically the solvent. remember that the solute is what there is less of and the solvent what there is most of.

OpenStudy (anonymous):

ok so when I'm now writing this out for homework how should i write it for my teacher to understand with out using a lot of words to describe it ?

OpenStudy (aaronq):

okay so you would first show how you found the mass using the density formula \(\rho=\dfrac{m}{V}\) then you could just use the formulas (as I'm sure they expect you to know what the symbols are) but do make sure you write units on everything.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

TTHANK YOU

OpenStudy (aaronq):

no problem